# tableReferences for Table S1

### v1

@@ 3 Comparison and scalability
These differences can be even more pronounced in applications where we only have a limited amount of data points, e.g., identifying rare cell types (Marouf et al., 2020), or where we have very high-dimensional data, e.g., neural population recordings in neuroscience (Stringer et al., 2019). In such cases, one needs to ensure that the distance measures can reliably distinguish different distributions for the given sample set size while remaining computationally tractable. In the following sections, we investigate the sensitivity of the three presented distances which do not rely on embeddings, when it comes to distinguishing data sets with varying numbers of samples (Section 3.1) and varying numbers of dimensions (Section 3.2). As the absolute values of the distance measures are often hard to interpret and different measures are on different scales, we examined the relative distances by comparing two or more models to the true data. We, therefore, applied the distances to compare samples from a ‘true’ distribution against itself (intra-dataset) and against samples from another distribution that is either an approximation or a shifted version of the true distribution (inter-dataset).
@@
In our experiments, we focused on inter- and intra-dataset comparisons for the following three datasets: First, we compared the two-dimensional Mixture of Gaussians (“2d-MoG”) dataset introduced in Fig. 1 with samples from a unimodal Gaussian approximation with the same mean and covariance as the true data samples. Second, we compared samples from a ten-dimensional standard normal distribution with a shifted normal distribution for which the first dimension is shifted by one (“10-dim Gaussian”). And last, we compared a standard normal distribution with varying dimensionality to shifted distributions for which we respectively shifted the first dimension for the inter-dataset comparison (“n-dim Gaussian”). We used default parameters for the SW and C2ST measures while adjusting the bandwidth parameter for the MMD measure for each of the three comparisons. Finally, as FID and other network-based distances do require an embedding network, we investigate the scaling properties specific to FID on the ImageNet dataset (in Section 3.3).
3.1 Varying number of samples
@@
We explored the robustness of the distances to low sample sizes on the 2d-MoG and the 10-dim Gaussian dataset. We found that for the 2d-MoG dataset, all measures failed to reflect the dissimilarity of the distributions at the lowest sample size of 50 samples (Fig. 8a). However, C2ST’s behavior differs from MMD and SW, with C2ST indicating that the distributions are similar (C2ST $\approx$ 0.5) for both intra- (true) and inter-dataset (approx.) comparisons while the other two distances indicate they are different (distance $\neq 0$). The malfunction of C2ST can be harder to detect in such cases, compared to the one of MMD and SW. While the latter is easily identified by the incorrect intra-dataset results, the malfunction of C2ST is hard to detect for unknown distributions. For all measures, computed values quickly stabilized by a sample size of 1000 and yielded the expected results of low intra-dataset differences and high inter-dataset differences.
@@
For the 10-dim Gaussian dataset, we observed that all distances can identify samples from the same distribution as more similar than samples from different distributions (Fig. 8b); while all three distances struggle with low sample sizes (see also Supp. Fig. S3, given enough samples, they all become robust, with no measure being clearly superior to the others. For the 2d-MoG experiment, more samples are required to clearly detect the difference between the two distributions (Fig. 8a) as compared to the case where the mean in one dimension is shifted (Fig. 8b). Intuitively, the larger the differences in the distributions we want to compare, the fewer samples we need to detect these differences (see additional experiments Supp. Fig. S3).
3.2 Varying dimensionality
@@
We further tested how the distances scale with the data dimension using $n$-dimensional standard normal distributions. In the first experiment, the shifted distribution differed only in the first dimension, which was mean-shifted by one. As the dimensionality increases, all distances reliably indicate no difference in intra-dataset comparisons , but C2ST is the only measure that consistently identifies the inter-dataset difference (Fig. 8c). Note that this dataset is different from Fig. 5b, where mean shifts were applied to every dimension. When we changed the structure of the data distribution (e.g., by changing the mean of all dimensions or their variances, see Supp. Fig. S4), we observed a similar picture with some particularities: While the SW distance with a fixed number of slices has difficulties if the disparity between the distribution is only in one dimension, its performance drastically improves for differences in all dimensions, which is expected from the random projections SW is performing. C2ST seems to robustly detect differences even in high dimensions in these modified datasets, though previous experiments showed that this measure can be oversensitive to small changes in high dimensions (Fig. 5b,c). Lastly, MMD is not robust across different dimensions for a Gaussian kernel with a fixed bandwidth. We could make the MMD robust across dimensions by using median heuristic (Section 2.3), where we would increase the chosen bandwidth such that it stays on the order of the euclidean distance between datapoints. Note that MMD can in general be highly sensitive to its hyperparameters, and an appropriate value depends not only on the dimensionality of the data (Supp. Fig. S5 and S6), but also on the structure of the distributions (Supp. Fig. S4).
@@ 3.3 FID-like distance comparison on ImageNet
Figure 9: Comparison of distances for ImageNet. (a) A comparison between the ImageNet test set and samples generated by an unconditional diffusion model, with varying sample sizes. (b) Distance evaluation on dog classes (D) versus non-dog classes (~D), highlighting differences in image representation between these two categories of real data. (c) Distances between sets of randomly selected images vs varying number of included classes of images (from 10 to 1,000) from the test set, using synthetic samples created by a conditional diffusion model.
@@
To also explore scaling properties of FID-like distances, we used images from the ImageNet dataset and embedded them with the InceptionV3 network (Deng et al., 2009, Szegedy et al., 2015b), following the implementation of Heusel et al. (2017). In addition to the 100,000 images in the ImageNet test dataset (1000 classes, 100 images per class), we generated high-quality synthetic samples using a state-of-the-art diffusion model as described by Dockhorn et al. (2022). We first produced 50,000 samples with the base unconditional version of this model. Using a conditional generative model, we additionally generated 100,000 class-conditional images (i.e., 100 per class), exactly matching the class distribution of the test-set. All of these images were embedded using the pre-trained InceptionV3 network (Szegedy et al., 2015a), transforming the raw images into a 2048-dimensional feature space.
@@
Calculating the FID involves computing the mean and covariance of the distributions in the embedding space and then calculating the squared Wasserstein distance analytically. However, we broadened our evaluation by applying additional distances to the distributions in the embedding space. While SW distance, C2ST, and FID effectively highlight the greater dissimilarity of synthetic samples to real images (i.e., the ImageNet test set) ( Fig. 9a) even for low sample sizes, the distinctiveness of the FID becomes only apparent when analyzing more than around 2000 samples. Estimating the full covariance matrix of the 2048-dim features with fewer samples leads to degeneracy and, thus, to numerical issues computing the square root. Common implementations hence generally recommend using more than 2048 samples (Heusel et al., 2018a; b). As also shown by Jayasumana et al. (2023), Betzalel et al. (2022), the Gaussian assumption in the FID is violated and can lead to problematic behavior. In contrast, the other distances reliably estimate a larger inter-dataset distance in regimes with few samples.
@@
In our subsequent analysis, we aimed to determine the effectiveness of various distances in discriminating between images from different classes by comparing the distances between different levels in the WordNet hierarchy (Miller, 1995). To this end, we focused on comparing images of dogs (D) with those of non-dog (~D) images (Fig. 9b). All investigated distances, except FID, were successful in identifying images from different classes as being more distinct than images from the same class. The FID comparison between data with multiple classes (~D vs. ~D) is higher than across dog classes and other classes (D vs. ~D). This shows that comparing two image classes can be problematic with FID, which is usually used to compare two distributions over many image classes (i.e. over natural images).
@@
To examine the effects only including a subset of classes from the dataset (i.e., modes of the distribution), we employed a test set and a conditional synthetic dataset, each comprising 1,000 classes with 100 samples per class. Our analysis involved comparing the complete test set against synthetic datasets that included only subsets of classes (Fig. 9c). For comparison purposes and as a control measure, we also conducted a scenario where, instead of selectively excluding classes, we randomly removed an equivalent number of images from the dataset. This approach revealed that limiting the dataset to a small number of classes compromised the performance across all evaluated distances, in contrast to the outcomes observed when randomly excluding a subset of images. To achieve performance comparable to that observed with random removals, it was necessary to include at least 800 classes in the comparison. As the InceptionV3 network is, in essence, trained to classify ImageNet images (Szegedy et al., 2015a) (under certain regularization schemes), hence the extracted high-level features may also be very sensitive to class-dependent image features and not necessarily for general image quality. This behavior can be observed in Fig. 9c and was recently explored by Kynkäänniemi et al. (2022). By replacing InceptionV3 with other embedding networks (e.g. CLIP which is trained to match images to captions) this class sensitivity can be reduced (Kynkäänniemi et al., 2022).
3.4 Computational complexity
While we only considered the sensitivity and specificity of the different distances in the previous paragraphs, we want to highlight that they differ also in their computational complexity. With respect to number of samples $N$, MMD and FID have a complexity of $O(N^{2})$ (note that for MMD the computational cost can be reduced, potentially at the cost of making approximation; Gretton et al. 2012a, Zhao & Meng 2015, Cheng & Xie 2021, Bodenham & Kawahara 2023, Bharti et al. 2023, Gretton et al. 2012b). While SW distance scales with $O(N\log N)$ (Nadjahi, 2021), it is difficult to make principled assessments of the computational complexity for C2ST, as it is highly dependent on the chosen classifier. But as more samples lead to larger training and test datasets, sample size is likely to influence the compute time. Similarly, the computational complexity of computing these distances increases as the dimensionality of the data increases, with non-trivial scaling depending on the task and hyperparameters chosen.
Note that despite their differences, all presented distances are reasonably tractable in the settings of our experiments, whereas the scaling experiments might be computationally unfeasible for other distances or datasets. We therefore strongly recommend carefully considering the complexity of the measure before conducting experiments on high-dimensional or very large datasets.
@@ 4 Scientific applications
To demonstrate how the presented distances apply to evaluating generative models of scientific applications, we focus here on two examples: decision modeling in cognitive neuroscience and medical imaging. For each application, we used two generative models or simulators to sample synthetic data. We then compared the synthetic samples to real data using the discussed distances. To obtain baseline values for each distance, we computed distances between subsets of real data. For SW distance, MMD, and FID we anticipated values proximal to zero, while for the C2ST, we expected a value around 0.5. These baseline assessments provide a lower threshold of model fidelity to which we compared the deviation of model-generated samples.
@@ 4.1 Models of primate decision making
Figure 10: Comparing models of primate decision making. (a) Schematic of a drift-diffusion model (DDM), a classical neuroscientific model of decision making behavior. Overall evidence drives the model toward one of two choices (drift), but sensory and environmental noise result in random fluctuations in evidence integration (diffusion). (b) Distributions of primate decision times from the real dataset (black), and two fitted models of varying complexity: DDM1 (gray) and DDM2 (gold). (c) SW distance, C2ST, MMD (bandwidth=0.5) between subsets of the three distributions. FID is not applicable in these comparisons, because the data are one-dimensional distributions. Scatter-points indicate comparisons between ten random subsets from each dataset. Thick horizontal bars indicate median values.
We explored the fidelity of two generative models in replicating primate decision times during a motion-discrimination task (Roitman & Shadlen, 2002). We evaluated two versions of a Drift-Diffusion Model (DDM; Fig. 10a) (Ratcliff, 1978), a frequently used model in cognitive neuroscience. The two versions differ with respect to the drift rate, which is the speed and direction at which evidence accumulates towards a decision, and the decision boundaries, which determine how much evidence is needed to make a decision. Specifically the first version (DDM1) uses a drift rate that varies linearly with position and time, and decision boundaries that decay exponentially over time whereas the second version (DDM2) uses a drift rate and decision boundaries that are constant over time (for details, see Section A.9). We fitted each model against empirical primate decision times with the use of the pyDDM toolbox (Shinn et al., 2020), generated one-dimensional synthetic datasets, and compared each dataset to the actual primate decision time distributions. While the resulting distributions of decision times are visually similar (Fig. 10b), the DDM-generated distributions DDM1 and DDM2 are noticeably broader compared to the more tightly clustered real decision times. Moreover, the DDM1 distribution appears more similar to the real distribution than that of DDM2, which is shifted towards the left. As expected, the DDM1 model more precisely mimics the real data distribution, as compared to the DDM2, across the median values of the SW distance, MMD, and C2ST distances (Fig. 10c). For C2ST DDM1 and the real data distribution are even indistinguishable, with median C2ST values around $0.5$. This suggests that SW and MMD provide a more nuanced differentiation between the models.

### v2

@@ 3 Comparing distances: sample size, dimensionality, and more
In this section, we empirically compare these distances along several dimensions, with additional results presented in the Appendices. When evaluating (or training) generative models, it is important to understand that different statistical distances pay attention to different features of the generated samples. A key aspect to consider is that differences between distances can be especially pronounced in applications where we only have a limited amount of data points, e.g., identifying rare cell types (Marouf et al., 2020), or where we have very high-dimensional data, e.g., neural population recordings in neuroscience (Stringer et al., 2019). In such cases, one needs to ensure that the distance measures can reliably distinguish different distributions for the given sample set size while remaining computationally tractable. We therefore perform a number of empirical studies here: First, we investigate SW, C2ST and MMD when it comes to distinguishing data sets given varying numbers of samples and data with varying dimensionality (Section 3.1). Second, we investigate the scaling properties of FID on the ImageNet dataset (Section 3.2), as well as considering the other distances when computed on the InceptionV3 image embeddings. We also summarize the theoretical sample and computational complexities of the four discussed distances in Section A.6.
Furthermore, different distances might weigh differently the importance of sample diversity, or mode coverage, versus how well the modes of the true distribution are captured. If models only capture one mode, or a subset of the modes, i.e., exhibit mode collapse, they might have learned to generate realistic but unvaried samples. We illustrated the trade-off between mode collapse and mode coverage by optimizing mis-specified (Fig. S2) and well-specified (Fig. S3) models using the different distances (see also Section A.7 and Theis et al. 2016).
Figure 8: Scalability of different statistical distances with sample size and dimensionality. (a,b) Comparison of sample sets with varying sample size (between 50 and 4k samples per set) of a ‘true’ distribution, either with a second set of samples of the same distribution or with a sample set from a model consisting of an approximated/shifted distribution. We show the mean and standard deviation over five runs of randomly sampled data. Note that the subscript for MMD distances (bottom) denotes the bandwidth of the Gaussian kernel used for a given dataset and we report the squared distance for MMD. (a) Distances for the 2d-MoG example shown in Fig. 1 compared to samples from a uni-modal Gaussian approximation with the same mean and covariance. (b) Distances for a ten-dimensional standard normal distribution, with a model whose first dimension is shifted by one. (c) Distances based on 10k samples from a standard normal distribution with varying dimensions (between 5 and 1000). As in (b), the first dimension of the model is shifted by one. We show the mean and standard deviation over five runs of randomly sampled data. One MMD bandwidth was selected for all n-dimensional datasets.
Note that the absolute values of the distance measures can be hard to interpret and different measures lie on different scales. For this reason, we compare the distance between a set of sample from the model and a set of samples from the true distribution to the distance between two sets of samples from the true distribution.
3.1 Varying number of samples and dimensionality
@@
Here, we assessed SW, C2ST and MMD using three datasets: First, we compared the two-dimensional Mixture of Gaussians (“2d-MoG”) dataset introduced in Fig. 1 with samples from a model consisting of a unimodal Gaussian approximation with the same mean and covariance as the true data. Second, we compared samples from a ten-dimensional standard normal distribution with a model that is also a normal distribution that matches the true distribution in all but the first dimension—where the mean is shifted by one (“10-dim Gaussian”). Last, we repeated the second experiment, now taking a normal distribution with varying dimensionality as true dataset, and a model who again matches in all dimensions, except for the shifted first dimension (“n-dim Gaussian”). We used default parameters for the SW and C2ST measures while adjusting the bandwidth parameter for the MMD measure for each of the three comparisons.
Sample size
@@
We explored the robustness of the distances to low sample sizes on the 2d-MoG and the 10-dim Gaussian dataset. We found that for the 2d-MoG dataset, computed values quickly stabilized for all measures by a sample size of 1000 and yielded the expected results of indicating low differences between two sets of samples from the true distribution (true), and a higher difference between samples from the true and model distributions (approx.). All measures failed to reflect the dissimilarity of the distributions at the lowest sample size of 50 samples (Fig. 8a). Here C2ST’s behavior differs from MMD and SW, with C2ST indicating that the distributions are similar (C2ST $\approx$ 0.5) while the other two distances indicate they are different (distance $\neq 0$).
@@
For the 10-dim Gaussian dataset, we observed that all three distances can robustly identify samples from the same distribution as more similar than samples from different distributions with no measure being clearly superior to the others (Fig. 8b). For the previous 2d-MoG experiment, more samples were required to clearly detect the difference between the two distributions (Fig. 8a) as compared to the 10-dim Gaussian (Fig. 8b). Intuitively, the more pronounced the differences in the distributions we compare, the fewer samples we need to detect these differences (see additional experiments in Supp. Fig. S4).
Dimensionality
@@
We further tested how the distances scale with the data dimension using the n-dim Gaussian dataset. As the dimensionality increases, all distances indicate no difference between two sets of samples from the true distribution, except C2ST which is the only measure that consistently identifies samples from the true and model as distinct (Fig. 8c). When we changed the structure of the data distribution (e.g., by changing the mean of all dimensions or their variances, see Supp. Fig. S5), we observed a similar picture with some particularities: While the SW distance with a fixed number of slices has difficulties if the disparity between the distribution is only in one dimension, its performance drastically improves for differences in all dimensions, which is expected from the random projections SW is performing. C2ST seems to robustly detect differences even in high dimensions in these modified datasets, though previous experiments showed that this measure can be oversensitive to small changes in high dimensions (Fig. 5b,c). Lastly, while here MMD is not robust across different dimensions for a Gaussian kernel with a fixed bandwidth, one typically chooses the bandwidth to be on the order of the distance between datapoints (e.g., using the median heuristic; Section 2.3). Note that in general, MMD can be highly sensitive to kernel- and hyperparameter choice; an appropriate setting depends not only on the dimensionality of the data (Supp. Fig. S6, S7, and S8), but also on the structure of the distributions (Supp. Fig. S5). SW distance, on the other hand, appeared robust to number of random projections used (Supp. Fig. S1 and A.4).
@@ 3.2 FID-like distance comparison on ImageNet
To explore scaling properties of FID-like distances, we generated high-quality synthetic samples using a state-of-the-art diffusion model as described by Dockhorn et al. (2022), which we compared to images from the ImageNet dataset, which contains 1000 classes, 100 images per class. All images were embedded using the pre-trained InceptionV3 network (Deng et al., 2009, Szegedy et al., 2016), following the implementation of Heusel et al. (2017), transforming the raw images into a 2048-dimensional feature space.
Figure 9: Comparison of distances for ImageNet. (a) A comparison between the ImageNet test set and samples generated by an unconditional diffusion model with varying sample sizes. (b) Distance evaluation on dog classes (D) versus non-dog classes (~D), highlighting differences in image representation between these two categories of real data. (c) Distances between sets of randomly selected images vs. varying number of included classes of images (from 10 to 1,000) from the test set, using synthetic samples created by a conditional diffusion model.
Sample size
@@
We first produced varying number of samples with the base unconditional version of the diffusion model and compared those to samples from Imagenet (Fig. 9a). Calculating the FID involves computing the mean and covariance of the distributions in the embedding space and then calculating the squared Wasserstein distance analytically. However, we broadened our evaluation by applying additional distances to the distributions in the embedding space. While SW distance, C2ST, and FID effectively highlight the greater dissimilarity of synthetic samples to real images (i.e., the ImageNet test set) even for low sample sizes, the distinctiveness of the FID only becomes apparent when analyzing more than 2000 samples. This is in line with common implementations which generally recommend using at least 2048 samples (Heusel et al. 2018a; b). Another reason for the limited performance could be that the Gaussian assumption in the FID can be violated (Jayasumana et al., 2023, Betzalel et al., 2022). In contrast, the other distances reliably estimate a larger inter-dataset distance in regimes with few samples.
Class discriminability
@@
We aimed to determine the effectiveness of various distances in discriminating between images from different classes. To this end, we focused on comparing ImageNet images of dogs (D) with ImageNet images of other (non-dog) classes (~D; Huang et al. 2021). All investigated distances, except FID, were successful in identifying images from different classes as being more distinct than images from the same class. The FID comparison between data with multiple classes (~D vs. ~D) is higher than across dog classes and other classes (D vs. ~D). This indicates that FID may be comparably insufficient in recognizing dissimilarities between distributions that have clear distinctions (whether across individual classes or in a one vs. all scenario).
Mode coverage
@@
To examine the effects when only considering a subset of classes from the dataset (i.e., modes of the distribution), we used the diffusion model to generate class-conditional images matching the classses of the ImageNet test set (Fig. 9c). Our analysis involved comparing the complete test set against model-generated datasets that included only subsets of classes. For comparison, we created a control dataset by randomly sampling from all classes (dashed curve in Fig. 9c). This approach revealed that limiting the dataset to a small number of classes compromised the performance across all evaluated distances, in contrast to the outcomes observed when randomly excluding a subset of images—indicating that all distances successfully detected the lack of mode coverage. To achieve performance comparable to that observed with random removals, it was necessary to include at least 800 classes in the comparison. As the InceptionV3 network is, in essence, trained to classify ImageNet images (Szegedy et al. 2015; under certain regularization schemes), the extracted high-level features may also be very sensitive to class-dependent image features and not necessarily for general image quality. This behavior can be observed in Fig. 9c and was recently explored by Kynkäänniemi et al. (2023). By replacing InceptionV3 with other embedding networks (e.g., CLIP, which is trained to match images to corresponding text), this class sensitivity could be reduced (Kynkäänniemi et al., 2023).
Additional generative models
We additionally generated images using a consistency model (CM) for unconditional image generation (Song et al., 2023), to investigate how these metrics perform when evaluated on images created by different generative models. This model was trained on ImageNet 64x64, similar to the GENIE model (Dockhorn et al., 2022). Moreover, we included the following models: BigGAN (Brock et al., 2019), ablated diffusion model (ADM; Dhariwal & Nichol 2021), Glide (Nichol et al., 2022), Vector Quantized Diffusion Model (VQDM; Gu et al. 2022), Wukong (Wukong, 2022), Stable diffusion 1.5 (SD1.5; Rombach et al. 2022b) and Midjourney (Midjourney 2022; details in Appendix A.13). We evaluated the metrics (and multiple commonly used variants of KID and C2ST) for each model against the ImageNet test set (see Table S4).
While there is some agreement between the metrics regarding which model generates images closest to the ImageNet test set, there are also differences in the relative ordering across different metrics. As expected, the most recent unconditional models trained directly on ImagenNet 64x64 performed best in our evaluation (GENIE, CS), better than the two other unconditional generative models (BigGAN, ADM). The other models are text-to-image and thus only prompted to generate images from specific ImageNet classes (GLIDE, VQDM, Wukong, SD1.5, Midjourney). Interestingly, the prompted models performed better than older unconditional models (BigGAN, ADM) most of the time. Recall that all we evaluate is the similarity to the ImageNet test set; prompted versions might produce images from the correct classes but might contain differences in style or appearance compared to actual images in ImageNet. Despite the demonstrated class-sensitivity (Fig. 9c), the InceptionV3 embeddings are thus also sensitive to different “styles” of natural images. Note that in this case, being closer to ImageNet does not necessarily mean generating better images (based on human perception), but rather creating images that are more ImageNet-like.
@@ 4 Scientific applications
To demonstrate how the presented distances apply to evaluating generative models of scientific applications, we focus here on two examples: decision modeling in cognitive neuroscience and medical imaging. For each application, we used two generative models or simulators to sample synthetic data. We then compared the synthetic samples to real data (hold-out test set) using the discussed distances. To obtain baseline values for each distance, we computed distances between subsets of real data. For SW distance, MMD, and FID we anticipated values proximal to zero, while for the C2ST, we expected a value around 0.5. These baseline assessments provide a lower threshold of model fidelity to which we compared the deviation of model-generated samples.
@@ 4.1 Models of primate decision making
Figure 10: Comparing models of primate decision making. (a) Schematic of a Drift-Diffusion model (DDM), a classical neuroscientific model of decision making behavior. Overall, evidence drives the model toward one of two choices (drift), but sensory and environmental noise result in random fluctuations in evidence integration (diffusion). (b) Distributions of primate decision times from the test set (black), and two fitted models of varying complexity: DDM1 (gray) and DDM2 (gold). (c) SW distance, C2ST, MMD (bandwidth=0.5) between subsets of the generated and real data distributions. FID is not applicable in these comparisons, because the data are one-dimensional distributions. Scatter-points indicate comparisons between ten random subsets from each dataset. Thick horizontal bars indicate median values.
We explored the fidelity of two generative models in replicating primate decision times during a motion-discrimination task (Roitman & Shadlen, 2002). We evaluated two versions of a Drift-Diffusion Model (DDM; Fig. 10a;  Ratcliff 1978), a frequently used model in cognitive neuroscience. The two versions differ with respect to the drift rate, which is the speed and direction at which evidence accumulates towards a decision, and the decision boundaries, which determine how much evidence is needed to make a decision. Specifically, the first version (DDM1) uses a drift rate that varies linearly with position and time and decision boundaries that decay exponentially over time, whereas the second version (DDM2) uses a drift rate and decision boundaries that are constant over time (for details, see Section A.14). We fitted each model to empirical primate decision times with the use of the pyDDM toolbox (Shinn et al., 2020), generated one-dimensional synthetic datasets, and compared each dataset to the actual primate decision time distributions. While the resulting distributions of decision times are visually similar (Fig. 10b), the DDM-generated distributions DDM1 and DDM2 are noticeably broader compared to the more tightly clustered real decision times. Moreover, the DDM1 distribution appears more similar to the real distribution than that of DDM2, which is shifted towards the left. As expected, the DDM1 model more precisely mimics the real data distribution, as compared to the DDM2, across the median values of the SW distance, MMD, and C2ST distances (Fig. 10c). For C2ST, DDM1 and the real data distribution are even indistinguishable, with median C2ST values around $0.5$. This suggests that SW and MMD provide a more nuanced differentiation between the models.